OpenStudy (lolamona):

hey if anyone willing to help me with some precal questions please send me a message . i need helpp please be my saviorrr

OpenStudy (sunnnystrong):

Sure! Just close this question and post a new one

OpenStudy (lolamona):

do you think we can message privatly

OpenStudy (sunnnystrong):

Hmm well I would rather you post the questions so other users can also comment and add to the discussion (also, medals ;P )

OpenStudy (lolamona):

lol i have attachments

OpenStudy (lolamona):

but ok

OpenStudy (sunnnystrong):

It is easier to attach files on questions anyways :D

OpenStudy (lolamona):

here is the first 2 questions

OpenStudy (sunnnystrong):

@lolamona Okay so if \[f(x)=x^2-4x-5\]& \[g(x)=x^2-25\] & they're asking about: \[\left( \frac{ f }{ g } \right)(x)=\frac{ f(x) }{ g(x) }\] Than: \[\left( \frac{ f(x) }{ g(x) } \right)=\frac{ (x-5)(x+1) }{ (x+5)(x-5)}\]

OpenStudy (sunnnystrong):

*Notice how i factored the top & bottom.. because a term will cancel @lolamona what cancels?

OpenStudy (lolamona):

o yes i see

OpenStudy (lolamona):

and then you make x+5 and x+1 equal to zero right

OpenStudy (lolamona):

well x-5 on both the top and bottom

OpenStudy (lolamona):

cancel out

OpenStudy (sunnnystrong):

yep which is interesting... anywho the domain are all real numbers *** except when the denominator is zero... so what do you think it is .... if \[\left( \frac{ f(x) }{ g(x) } \right)=\frac{ (x-5)(x+1) }{ (x-5)(x+5) }\]

OpenStudy (sunnnystrong):

Second one --->

OpenStudy (lolamona):

cant equal cant equal -1 and 5 or am i wrong

OpenStudy (sunnnystrong):

no it can! heres why

OpenStudy (sunnnystrong):

sooo.. do you recall that you can't divide by zero?

OpenStudy (lolamona):

x-5 cause it cancels out

OpenStudy (sunnnystrong):

basically it's against the law in math lol & yeah i didn't mean to confuse you! if they were asking for the most simplified version of f/g(x) than yeah... the x-5 would cancel out. BUT it is still in the domain

OpenStudy (sunnnystrong):

so the domain is all real numbers EXCEPT 5 & -5

OpenStudy (sunnnystrong):

aka ZEROS of the denominator

OpenStudy (lolamona):

ohhhhhhhh lol

OpenStudy (sunnnystrong):

Yep XD now the next one

OpenStudy (sunnnystrong):

Recall that: \[(f og )(x)=f(g(x))\] Sooo basically this means f composed with g... If: \[g(x)=\frac{ 1 }{ x-1 }\] \[f(x)=x^2+6x+5\] Than: \[f(g(x))=(g(x))^2+6(g(x))+5\]

OpenStudy (lolamona):

so plugging in f and g

OpenStudy (sunnnystrong):

Yep would look something like this... \[f(g(x))=\left( \frac{ 1 }{ x-1 } \right)^2+6\left( \frac{ 1 }{ x-1 } \right)+5\]

OpenStudy (lolamona):

ok let me write this down lol

OpenStudy (sunnnystrong):

simplify--> oh shoot messed up a sign ah

OpenStudy (sunnnystrong):

x+1**

OpenStudy (sunnnystrong):

\[f(g(x))=\left( \frac{ 1 }{ x+1 } \right)^2+6\left( \frac{ 1 }{ x+1 } \right)+5\] \[f(g(x))=\left( \frac{ 1 }{ x+1 } \right)\left( \frac{ 1 }{ x+1 } \right)+\left( \frac{ 6 }{ x+1 } \right)+5\] \[f(g(x))=\left( \frac{ 1 }{ x^2+2x+1 } \right)+\left( \frac{ 6 }{ x+1 } \right)+5\]

OpenStudy (sunnnystrong):

& alternate forms of that ^ \[f(g(x))=\frac{ 5x^2+16x+12 }{ x^2+2x+1 }\]

OpenStudy (lolamona):

yay i got that right

OpenStudy (sunnnystrong):

Woo :D good job!

OpenStudy (lolamona):

lol think you can help me with a couple more you explain very well

OpenStudy (sunnnystrong):

yeah sure i would love too :D

OpenStudy (sunnnystrong):

Okay so the first one--> we are working with an arithmetic sequence: If the sequence is 18,54,90,126 *Notice that each therm is increasing by 36. If: \[a_{n}=a_{n-1}+d\] & an= nth term in sequence an-1=the term before the nth term in the sequence & d=the common difference than the formula is : \[a_{n}=a_{n-1}+36\]

OpenStudy (sunnnystrong):

So: they're giving us up to the 4th term in the sequence & they're asking about the eighth term: Basically there's the logic ^ But you can just multiply 36*4 and add it the the 4th term and you arrive at the same results

OpenStudy (sunnnystrong):

@lolamona sooo... what is the eighth term in the sequence? aka the distance the projectile traveled in 8 seconds

OpenStudy (lolamona):

hold on

OpenStudy (lolamona):

im lost -.-

OpenStudy (sunnnystrong):

Hmm.. okay so lets work with the arithmetic sequence actually

OpenStudy (3mar):

With @sunnnystrong, you are in good hands, @lolamona ....

OpenStudy (sunnnystrong):

\[a_{n}=a_{n-1}+36\] \[a_{1}=18\] \[a_{2}=18+(2-1)36\] \[a_{2}=18+(1)36=54\] \[a_{3}=18+(3-1)36=90\] \[a_{4}=18+(4-1)36=126\] \[a_{5}=18+(5-1)36=162\] \[a_{6}=18+(6-1)36=194\] \[a_{7}=18+(7-1)36=226\] \[a_{8}=18+(8-1)36=258\]

OpenStudy (sunnnystrong):

@lolamona there's the sequence.. up to the eighth term.\[a_{n}=a_{1}+(n-1)d\] d= common difference (32) a1=18 (aka first term in sequence)

OpenStudy (lolamona):

so 18 is the first number so a1 and what ever number you are trying to find you put subtracted by 1 and the common difference is 32

OpenStudy (sunnnystrong):

Yep :D

OpenStudy (sunnnystrong):

okay i messed up somewhere ^^ one sec haha

OpenStudy (lolamona):

for 8 its 270

OpenStudy (sunnnystrong):

from the 6th term --> guess i messed up simple math ok lol a6=198 a7=234 a8=270 okay lets do the next :D

OpenStudy (lolamona):

kk

OpenStudy (sunnnystrong):

\[s_{n}=\frac{ 1 }{ 2 }(a_{1}+a_{n})\] sn= the sume of the first n #'s inn the sequence (in this case up to 24 & prior) a1= first # in secuence an= nth number in sequence

OpenStudy (sunnnystrong):

\[s_{24}=\frac{ 1 }{ 2 }(a_{1}+a_{24})\]

OpenStudy (lolamona):

we will have to find a1 no

OpenStudy (sunnnystrong):

pretty much yea :P

OpenStudy (sunnnystrong):

so if we know the common difference (aka d) is 3.7 & \[a_{13}=1.9\] than: \[a_{13} =a_{1}+12d=1.9\] Solve for a1 a1=-42.5 find a24: \[a_{24} =-42.5+(24−1)3.7=42.6]

OpenStudy (lolamona):

omg thank you so much for explaining it so well i really appreciate your help and patience

OpenStudy (lolamona):

can i message you anytime i need help with precalc you have really helped me alot

OpenStudy (sunnnystrong):

no problem! :D always happy to help! so a24=42.6 a1= -42.5 Than--> \[S_{24}=.5(24)((-42.5)+(42.6))=1.2\]

OpenStudy (sunnnystrong):

phewwww finally done with that sequence stuff :D good job & yeah sure if i can i will try to help :)

OpenStudy (lolamona):

lol thank you , you have been an amazing help

OpenStudy (sunnnystrong):

thanks you :) !